# Fast Real Evaluation Through Sound Mixed-Precision Tuning

### v1

@@ 6.2. Evaluation
Reval evaluates each instruction in turn, at that instruction’s precision assignment. Then, for each output register, Reval uses the conversion and distance functions to determine if the computed intervals are sufficiently narrow. If not, Reval performs a precision assignment using the exponent trick formulas and the already-computed intervals for each register, and repeats this process until a sufficiently-narrow interval is found. This minimal evaluation loop reduces evaluation overhead, and performance is largely determined by the choice of precision for each instruction.
If the precision assigned to a specific instruction does not change (or even decreases) from one iteration to the next, there’s no need to recompute that instruction, as long as its inputs are not recomputed. Reval checks for this condition during precision assignment and sets a bit in a per-instruction “repeat vector”; instructions with the repeat bit set are not re-executed. On the other hand, if an operation is ever assigned a precision higher than a user-specified precision bound, Reval terminates execution and signals an error. This “early stopping” can trigger as early as the first iteration and saves significant time on points Reval cannot evaluate. Reval can also exit the evaluation loop if the output intervals’ error flag is set to $\top$, if movability flags prove the expression unevaluable, or if a user-defined iteration limit is reached.
6.3. Precision Tuning
Reval’s precision tuning pass is basically Equation 3 converted to code. Reval allocates a vector $P[z_{i}]$ and fills the vector starting from the end. The already-filled, later elements of the array represent the precisions $p_{i}$ for each instruction, while the initial, un-filled elements represent a target precision for $\operatorname{bound}(\hat{z}_{i})$. In code, the algorithm is:
@@ 6.3. Precision Tuning
The $\operatorname{intro}$ and $\operatorname{ampl}$ terms are computed using the exponent trick formulas of Tables 2 and 2. This single “backward pass” over the precision array suffices to compute a precision assignment for each instruction. Reval performs some additional, minor optimizations to make this formula even faster to compute, including pre-propagating the “${}+2$” factors and computing the $\operatorname{intro}$ and $\operatorname{ampl}$ terms simultaneously.
Soundness ensures that Reval terminates: without a sound precision assignment, it would be possible to enter an infinite loop, where at each iteration the same precisions are assigned. Reval does include code to check for this condition (in case the user provides a custom discretization with a too-low target precision $t$) but this code is not triggered in our evaluation. However, achieving termination also requires some additional pragmatic changes to the simple theory of Equation 3.
6.4. Slack
@@
Sometimes, an interval crosses zero, in which case $\operatorname{minlog}(\bar{z})$ is not defined—it approaches negative infinity. Similarly, some intervals are unbounded, in which case $\operatorname{maxlog}(\bar{z})$ approaches positive infinity. In both cases, the exponent trick formulas would evaluate $\operatorname{intro}$ and $\operatorname{ampl}_{k}$ as infinity, which of course is not a usable precision assignment. Intuitively, what’s happening is that the precision assignment algorithm does not know how many bits will suffice to eliminate, say, a catastrophic cancellation, and assigns infinitely many bits to remain sound.
In this case, Reval falls back to Ziv onion peeling; this is why Reval can require multiple iterations of precision assignment and evaluation before finding a narrow-enough interval. Specifically, Reval uses $\operatorname{slack}[n]$ in place of any infinity in the exponent trick formulas, where $n$ is how many iterations Reval has performed and $\operatorname{slack}[n]$ is a guess as to how many bits might be needed. So, for example, the interval $\bar{z}=[-2^{-93},2^{-108}]$ has $\operatorname{minlog}(\bar{z})=\min(-93,-108)+\operatorname{slack}[n]$. Slack is equal to 512 after the first iteration, and doubles for every further iteration.
This slack mechanism reintroduces a naive precision-doubling algorithm, but only partially:
  • The precisions are still not uniform;
  • The non-zero, non-infinite endpoints still provide useful information about the necessary precision, so even though slack terms involve guesses they are at least somewhat informed; and
  • Slack typically affects only one or a few operations, and the other operations still set their precisions using sound mechanisms.
Because slack only affects a few operations, it can be set more aggressively. The initial guess of 512, for example, is much larger than Sollya’s initial guess of 64.
6.5. Correct rounding
Reval also uses the slack mechanism to achieve correct rounding. Correct rounding requires firmly placing the value to one side or another of a rounding boundary in the target precision. The challenge is that the output can be very close to such a rounding boundary, so that proving it to be on one side or another can require $\bar{z}$ to be much narrower than merely $2^{-p}$. Note that this is rare: in most cases, the two endpoints are either more than 1 ULP apart or match exactly; however, correct rounding is valuable for a variety of use cases (Brisebarre et al., 2024) due to its determinism.
Reval uses the slack mechanism to handle this case: if the output interval rounds to two successive floating-point numbers in the target precision, meaning that Reval has achieved a faithfully- but not correctly-rounded result, the target precision $t$ is increased by $\operatorname{slack}[n]$ before performing precision assignment.
@@ 6.6. Other pragmatic tweaks
Rival, the interval library Reval uses, does not seem to provide operators that are tight in the sense of IEEE 1788 (iee, 2015). This means that Equation 3 can produces slightly-too-low precision assignments. In practice, this is not much of a problem, because the interval operators that aren’t tight, like $\sin$ and $\cos$, also use over-approximating exponent trick formulas. But as extra security to avoid issues, Reval uses “${}+5$” instead of “${}+2$” in its precision assignment formulas. This allows for up to 3 ULPs of looseness in each operator, which we believe is more than Rival needs. [^14] We’ve validated this with some experiments, but to our knowledge Rival has not been proven correct and verified bounds are not available. Since much of runtime is spent at very high precisions, the addition of a few bits doesn’t impact performance much.
Reval also supports a number of operators that have no exponent trick formula in Table 2, such as $\operatorname{erfc}$. For these operators Reval simply uses $\operatorname{slack}$ for their $\operatorname{ampl}$ factors. Naturally, Reval’s performance is then no better than naive precision-doubling, but these operators are rare so basic support seems acceptable. For some operators, the slack mechanism is only resorted to for some inputs. For example, $\lceil\log_{2}\operatorname{ampl}_{1}(\operatorname{asin}(\hat{x}_{i}))\rceil\leq\operatorname{maxlog}(\bar{x}_{i})-\operatorname{minlog}(\bar{z}_{i})+1$ as long as $\operatorname{maxlog}(\bar{z}_{i})\leq 0$, but when that condition is not true, slack still has to be added. Similarly, for the rounding operators like $z=\lfloor x\rfloor$, $\lceil\log_{2}\operatorname{ampl}_{z}(\hat{z})\rceil$ is set to $\operatorname{maxlog}(\bar{x})$ if $\hat{z}$ contains more than one integer, $\infty$- ∞ if it contains zero integers, and $\operatorname{slack}[n]$ if it contains exactly one integer (meaning the true value of $x$ might be right near a rounding boundary).
Finally, the $\operatorname{logspan}$ terms in the exponent trick formulas caused issues in early iterations when intervals are very wide. We therefore set $\operatorname{logspan}(\bar{z})=0$; surprisingly, this does not affect soundness, because in the final iteration, where the slack mechanism isn’t used, all values must be known to at least 2 bits of precision, meaning $\operatorname{logspan}(\bar{z})\leq\frac{1}{4}$. Since no exponent trick formula contains more than three $\operatorname{logspan}$ terms, the total added precision from $\operatorname{logspan}$ terms is less then 1. Reval’s use of “${}+5$” instead of “${}+2$” covers this extra bit as well.
@@ 1. RQ1
    Can Reval evaluate mathematical expressions faster than a uniform-precision baseline?
@@ 2. RQ2
    Can Reval evaluate more mathematical expressions than a uniform-precision baseline?
@@ 3. RQ3
    Does Reval use lower precisions, on average, than a uniform-precision baseline?
We evaluate all three research questions on a set of 528 benchmarks drawn from the Herbie 2.0 benchmark suite, each evaluated at 8256 inputs. For the uniform-precision baseline, we use both Reval with precision tuning disabled and also the state-of-the-art Sollya tool (except for RQ3).151515 Other tools that perform real evaluation include Wolfram Mathematica (Wolfram, 2020), $\lambda_{s}$ (Sherman et al., 2021), the Android and Windows calculators (Boehm, 2020), Calcium (Johansson, [n. d.]), and possibly others. Prior work (Flatt and Panchekha, 2023) has found Mathematica to be much slower than Reval, and the other libraries are not, to our knowledge, optimized for performance.
@@ 7.1. Methodology
The 528 benchmarks are mathematical expressions written in the FPCore 2.0 format (Damouche et al., 2016). Benchmarks are short, with 1–59 operations (average 11.6) over 1–9 variables (average 2.5). The most common operations by far are the arithmetic operations, with multiplication at over 30% of operations and addition, subtraction, and division constituting another 50% of operations. The most common non-arithmetic operations are the exponential, power, sine, cosine, logarithm, and square root operations, each approximately 1.5% of operations. The rarest operations were arctangent, hypotenuse, and fused multiply-add. For each benchmark, we randomly sample 8256 input points (that is, values for each variable) using the Herbie 2.0 tool and use the same input points for Reval and both baselines. Input points are sampled using rejection sampling, with inputs that have a domain error according to the baselines or Reval being rejected.
We discard benchmarks with loops, conditionals, data structures (arrays or tensors), or functions not supported by Sollya (mainly cbrt and atan2), which eliminates 17 of the full 547-strong benchmark suite. In addition, two benchmarks (that we know of) trigger a bug in Sollya related to the subexpressions $z^{z}e^{-z}$ and $e^{-z}\ell^{\exp(z)}$, reported to and acknowledged by the developers. We remove these two benchmarks as well. 528 remain, of which 29 are evaluated to single-precision accuracy, while 499 are evaluated to double-precision accuracy; inputs are also given in the same precision. Sollya doesn’t provide hypot and fma functions so these are expanded into $\sqrt{x^{2}+y^{2}}$ and $x\cdot y+z$ during compilation to Sollya’s input language. 98 benchmarks specify a precondition, some of which are simple ranges but others of which involve complex conditionals. Since Sollya does not support preconditions, we do not time precondition evaluation, but use only input points that satisfy the preconditions. Reval is set to correctly round each expression on each of its inputs; however, Sollya instead only attempts to round faithfully.161616 Sollya’s definition of faithful rounding allows it to round a value to any of three possible floating-point values, presumably in the case of exact floating-point results. This is rare but does happen in our evaluation. Moreover, in some cases, the correctly-rounded result is infinite (for example, if the real result is $10^{400}$). In this case, Reval returns infinity but Sollya’s results vary; we track these input points separately and ignore them in RQ1–3.
@@
All tools are run with Sollya’s default minimum and maximum precisions or $10+t$ and $(10+t)\cdot 512$, where $t$ is either 24 or 53. These initial and maximum precisions are Sollya’s defaults. For Sollya the necessary precision indicated by setting the prec variable. Sollya’s main evaluation strategy is to evaluate the whole expression under a uniform precision with interval arithmetic, doubling the precision until a faithful rounding is found. However, Sollya does, in various special cases, apply additional tricks, including (from an examination of the source code) Taylor series, l’Hôpital’s rule, and in some cases smaller precision changes than doubling. To the best of our knowledge, however, it does not perform mixed-precision evaluation, except in the case of constants that are cached at high precision. Reval applies the precision tuning algorithm described in this paper, with failure after 5 iterations. We also compare to “Baseline”, which is Reval but applying a uniform precision that doubles with each iteration until the correct rounding has been obtained or the maximum precision has been reached.
For each tool, benchmarks are compiled from FPCore to the tool’s input syntax. The compilations are all straightforward. For Sollya each expression is first defined as a function, with the compilation time not measured, and then that function is evaluated on each input point in turn. The separate compilation step allows Sollya to perform compile-time optimization or to share caches across inputs. Reval also performs a separate compilation step; however, it does not perform any caching except for exact constants like 1. To check that the compilation is performed correctly, we compare the outputs from all three tools (which is how we discovered the previously-mentioned Sollya bug).
@@
Evaluation time for an input point must be measured carefully. Reval and Baseline are called within the same process, so a simple nanosecond timer is used. Sollya, however, is started as an external process. To avoid including parsing time and other overhead, Sollya’s time command is used to measure its running time. All tools are subjected to a 20 millisecond timeout. Sollya in particular suffers from what seem like infinite loops; 20 milliseconds was selected as the shortest timeout that still allows Sollya to successfully evaluate most of even the most difficult inputs. A larger timeout would only bias the results toward Reval. Evaluation is performed on a machine with an Intel i7-8700 at 3.7 GHz with 32 GB of DDR4 memory running: Ubuntu 22.04; Sollya 7.0 with MPFI 1.5.3, MPFR 4.1.0, and GMP 6.2.1; and Reval 2.0 with Racket 8.11.1. The evaluation code itself is single-threaded, as are Reval and Baseline and, to our knowledge, Sollya, and the evaluation is run without load on the machine.
@@ 7.2. Results
Figure 3 shows our results for RQ1. The figure shows evaluations per second for Reval and Baseline, relative to Sollya. Only the 4,359,168 input points where all three tools successfully evaluate the expression are included.171717 This way the handling of invalid inputs, unsamplable inputs, precision limits, and so on do not influence the results. Input points are aggregated by “difficulty”, which is the number of iterations Reval needs to evaluate the point; higher difficulty typically means higher precision. Note that this difficulty metric (chosen because it is easy to compute) is biased against Reval: a point that takes Reval many iterations may nonetheless be easy for Sollya or Baseline. Very roughly, every difficulty level has about 20×20$\times$20 × fewer points than the previous; there are only 30 difficulty-5 points, so its timing results are somewhat unreliable, the the relative standings of the three tools occur repeatedly.
Reval achieves significantly more evaluations per second, both on average and at any difficulty level. Averaged over all points, Reval is 1.25×1.25$\times$1.25 × faster than Sollya and 1.21×1.21$\times$1.21 × faster than Baseline. In fact, Baseline and Sollya perform at similar speeds across the difficulty spectrum, as expected given their similar precision assignment strategy. Reval has the smallest advantage at difficulty 2, which we believe is due to the slack mechanism sometimes overestimating the necessary precision. Reval’s advantage grows as difficulty increases: for the most difficult points, Reval is 2.99×2.$9$ faster than Sollya and 1.71×1.71$\times$1.71 × faster than Baseline. This is expected: more difficult points require higher precisions meaning that precision assignment provides a greater benefit.
Figure 2. Reval’s, Sollya’s, and Baseline’s speed by difficulty of input point. Sollya is normalized to 1.0; higher is better. All three tools perform similarly on the easiest points, and Baseline and Sollya perform similarly throughout, as expected given their very similar core algorithms. However, Reval has a growing advantage on harder inputs, with a speed-up over Sollya of 1.25×1.25$\times$1.25 × overall and 2.99×2.$9$ on the most difficult inputs.
Figure 3. The number of points evaluated by Reval, Sollya, and Baseline, broken down by whether the correctly-rounded result was zero or a non-zero real number. Higher is better. Reval and Baseline both evaluate far more points than Sollya. Reval and Baseline are largely comparable, showing that Reval’s precision tuning does not much affect whether points are successfully evaluated.
@@
Figure 3 shows our results for RQ2. This graph considers only points successfully evaluated by at least one, but not all, of the tools. Reval and Baseline evaluate a similar number, with Reval evaluating more points where the correct answer is 0 (3743 to Baseline’s 3741) but fewer points where the correct answer is non-zero (10 to Baseline’s 54). This is expected from precision tuning: Reval can sample more points by avoiding timeouts, but its overapproximations can also cause it to unnecessarily pass the precision limit. In any case, the differences are minimal, showing that precision tuning’s throughput gains come more or less “for free”. Sollya, meanwhile, evaluates only 1635 zero values and 52 non-zero values. We believe the gap is due to Reval and Baseline’s Rival interval arithmetic library handling overflow and underflow in a more rigorous manner (Flatt and Panchekha, 2023) than Sollya’s MPFI. Reval and Baseline also sample 100,346 more points where the correctly-rounded result is infinite; we believe this to be an intentional choice by the Sollya developers, so out of fairness they are not shown in Figure 3.
@@
Figure 4 shows our results for RQ3. This plot shows the distribution of execution time by precision across all operations, for both Reval and Baseline. (We were not able to log such granular runtime information in Sollya.) Reval saves a significantly fraction of runtime by executing operations at lower precision than Baseline—note the higher bar at precision $2^{6}$ in exchange for lower bars at precisions $2^{7}$ through $2^{11}$. Precision assignment is also fast, as shown by the separate column on the right: precision tuning consumes 4.1% of run time but saves 17% of run time. Reval spends a tiny fraction of runtime at high precision; in this sense, its precision assignments can be said to be close to optimal.
Figure 4. A weighted histogram showing the distribution of operation execution time by precision, for both Reval and Baseline. Lower is better. Baseline wastes a substantial fraction of its execution time re-computing the same operations at successively higher precisions, leading to a lot of wasted time at precisions between $2^{7}$ and $2^{11}$. By contrast, Reval spends much less time at these high precisions because it can assign high precision only to the few operations that need it.

### v2

@@ 6.2. Evaluation
Each instruction in the instruction sequence has its own precision, initialized to a constant. To evaluate the compiled expression, Reval repeatedly evaluates each instruction and then tunes the precision assignment, stopping once each output register holds an interval with equal endpoints under the discretization’s distance function. This minimal evaluation loop reduces evaluation overhead, and performance is largely determined by the choice of precision for each instruction.
Reval’s precision tuning pass is basically Equation 3 converted to code. Instruction precisions are stored in a vector $P[z_{i}]$ which is filled in a “backward pass” starting from the end. The already-filled, later elements of the array represent the precisions $p_{i}$ for each instruction, while the initial, un-filled elements represent a target precision for $\operatorname{bound}(\hat{z}_{i})$. In code, the algorithm is:
@@
The $\operatorname{intro}$ and $\operatorname{ampl}$ terms are computed using the exponent trick formulas of Tables 2 and 2. Reval performs some additional, minor optimizations to make this formula even faster to compute, including pre-propagating the “${}+2$” factors and computing the $\operatorname{intro}$ and $\operatorname{ampl}$ terms simultaneously.
If the precision assigned to a specific instruction does not change (or even decreases) from one iteration to the next, there’s no need to recompute that instruction, as long as its inputs are not recomputed. Reval checks for this condition during precision assignment and sets an instruction-specific bit in a “repeat vector”; instructions with the repeat bit set are not re-executed. This “repeat optimization” is only possible thanks to the non-uniform precisions; with a uniform precision, every instruction’s precision would increase at every iteration and no repeat bits would be set.
6.3. Slack and Correct Rounding
@@
Since Reval’s precision assignment algorithm is sound, the precisions assigned in the precision tuning pass are high enough that the next iteration will yield narrow-enough intervals produce a faithfully-rounded result. However, if an intermediate interval $\bar{z}$ either crosses zero or contains infinity, its $\operatorname{minlog}(\bar{z})$ or $\operatorname{maxlog}(\bar{z})$ approach infinity, which makes precision assignment impossible. Intuitively, what’s happening is that the precision assignment algorithm does not know how many bits will suffice to eliminate, say, a catastrophic cancellation, and assigns infinitely many bits to remain sound.
Reval addresses this issue with a partial return to precision doubling, replacing the infinite value with a guess. Specifically, Reval uses $\operatorname{slack}[n]$ in place of any infinity in the exponent trick formulas, where $n$ is how many iterations Reval has performed. So, for example, the interval $\bar{z}=[-2^{-93},2^{-108}]$ has $\operatorname{minlog}(\bar{z})=\min(-93,-108)+\operatorname{slack}[n]$. While this “slack mechanism” is similar in some ways to the standard precision-doubling algorithm, it has some key advantages. Precisions are still not uniform, and the non-zero, non-infinite endpoints of intervals still provide useful information to precision tuning. Moreover, slack typically affects only one or a few operations, and all other operations still set their precisions using sound mechanisms. Because of this last point, it can be set more aggressively: 512 for $\operatorname{slack}[0]$ and doubling in later iterations.
Even a precision assignment with no slack still only guarantees a faithful rounding, not a correct rounding, because even a very narrow output interval can still straddle a rounding boundary. Reval detects this case using the “neighbor” distance value and increases the target precision until the output interval falls to one side of the rounding boundary or the other. Since there’s no principled way to know how close to the rounding boundary the true value lies, Reval increases the target precision by $\operatorname{slack}[n]$. Users can disable this mechanism by providing a custom discretization which never returns the “neighbor” distance value, but it is on by default because correct rounding is valuable for a variety of use cases (Brisebarre et al., 2024).
6.4. Invalid Inputs
Inputs to a real expressions can cause domain errors (detected by Rival’s error flags) or simply be impossible to evaluate under a given precision bound. Reval thus exits the evaluation loop if:
  • an output register’s error flag is set;
  • an output register’s movability flags prove the expression unevaluable at any precision;
  • an operation is assigned a precision beyond a user-specified precision bound. [^13] Reval also supports a user-set iteration bound.
The last case, “early exit”, can be triggered as early as the first tuning iteration and is critical for Reval’s robustness to invalid inputs. That said, it can trigger inappropriately: the precision assignment over-estimates the necessary precision, so a lower precision, possibly lower than the precision bound, might still work. This edge case is rare but does occur in our evaluation.
To help address this issue, Reval has an optional “lower bound mode”. In this mode, a second set of exponent tricks formulas, which under-approximate the amplification factors, are used for the early exit calculation. This second set of formulas largely with switches $\operatorname{maxlog}$ for $\operatorname{minlog}$ and vice versa, with adjustments to the approximations used for some operations. If the precision assignment is larger than the precision bound but the lower-bound formula is below the precision bound, the precision assignment is set to exactly the precision bound. This avoids all cases of inappropriate early stopping in our evaluation, but also increases the time spent on unevaluable inputs, so this “lower bound mode” is off by default.
@@ 6.5. Other pragmatic tweaks
Equation 3 assumes the interval library produces maximally tight intervals (iee, 2015), but this is not actually true for Rival. As extra security to avoid issues, Reval thus uses “${}+5$” instead of “${}+2$” in its precision assignment formulas. This allows for up to 3 ULPs of looseness in each operator, which we believe is more than Rival needs. [^14] We’ve validated this empirically but Rival has not been proven correct and verified bounds are not available. Since much of runtime is spent at very high precisions, the addition of a few bits doesn’t impact performance much.
Reval also supports a number of operators that have no exponent trick formula in Table 2, such as $\operatorname{erfc}$. For these operators Reval simply uses $\operatorname{slack}$ for their $\operatorname{ampl}$ factors. Reval is then no better than standard precision-double algorithms, but these operators are rare so basic support seems acceptable. Alternatively, some operators like $\operatorname{asin}$ use the slack mechanism only for certain inputs. That is, $\lceil\log_{2}\operatorname{ampl}_{1}(\operatorname{asin}(\hat{x}_{i}))\rceil\leq\operatorname{maxlog}(\bar{x}_{i})-\operatorname{minlog}(\bar{z}_{i})+1$ when $\operatorname{maxlog}(\bar{z}_{i})\leq 0$, but when that condition is not true, slack is used. Similarly, for the floor operator, $\lceil\log_{2}\operatorname{ampl}_{z}(\lfloor\hat{x}\rfloor)\rceil=\operatorname{maxlog}(\bar{x})$ if $\hat{x}$ contains more than one integer, $\infty$- ∞ if it contains zero integers, and $\operatorname{slack}[n]$ if it contains exactly one integer.
Finally, the $\operatorname{logspan}$ terms in the exponent trick formulas are very large in early iterations; we therefore replace $\operatorname{logspan}(\bar{z})\mapsto 0$. Surprisingly, this does not affect soundness, because in the final iteration with no slack mechanism, all values are known to at least 2 bits of precision, meaning $\operatorname{logspan}(\bar{z})\leq\frac{1}{4}$. Since no exponent trick formula contains more than three $\operatorname{logspan}$ terms, the total added precision from $\operatorname{logspan}$ terms is less then 1, which is covered by Reval’s use of “${}+5$” instead of “${}+2$”.
@@ 1. RQ1
    Can Reval evaluate expressions faster than a uniform-precision baseline?
@@ 2. RQ2
    Is this because Reval uses lower precisions than a uniform-precision baseline?
@@ 3. RQ3
    How do Reval’s precision assignments contribute to performance improvements?
We evaluate all three research questions by comparing Reval to a variant, Baseline, that uses a standard uniform precision-doubling algorithm, on 567,238 inputs to 469 benchmarks from the Herbie 2.0 benchmark suite. For RQ1, we also compare to the state-of-the-art Sollya tool.151515 Wolfram Mathematica (Wolfram, 2020), the $\lambda_{s}$ (Sherman et al., 2021) and Marshall languages (Sherman et al., 2019), the CGAL and LEDA graphics packages (Seel, 2024; Näher and Mehlhorn, 1992), the Calcium (Johansson, [n. d.]) library, and the Android and Windows calculators (Boehm, 2020) also perform real evaluation but aren’t optimized for performance and use the similar uniform precision algorithms to Sollya and Baseline.
@@ 7.1. Methodology
Benchmarks
The Herbie 2.0 suite contains 509 mathematical expressions written in the FPCore 2.0 format (Damouche et al., 2016). Benchmarks are short, with 1–59 operations (average 11.6) over 1–9 variables (average 2.5). The most common operations by far are the arithmetic operations (80%+) but the exponential, power, sine, cosine, logarithm, and square root operations are also common (each $\approx 1.5\%$). We discard 14 benchmarks with unsupported operations [^16] Loops, conditionals, data structures (arrays or tensors), and the cbrt and atan2 functions not supported by Sollya. as well as two benchmarks that trigger a bug in Sollya, [^17] Sollya incorrectly handles underflow $z^{z}e^{-z}$ and $e^{-z}\ell^{\exp(z)}$; the bug was reported and acknowledged by the developers. leaving 493 total.
Inputs
For each benchmark, we randomly sample 8256 double-precision input points (that is, values for each variable) using the Herbie 2.0 tool [^18] Herbie 2.0 uses rejection sampling and samples points with no domain errors and which satisfy any benchmark-specific preconditions. and use the same input points for Reval and the baselines. All inputs to the remaining 493 benchmarks are evaluated to correctly-rounded double-precision values;191919 Sollya mostly produces correctly-rounded results, but in 5355 cases, roughly 1%, instead produces a faithful rounding. inputs whose correctly-rounded result is infinite are discarded because Sollya’s handling of these points is inconsistent. Of the remaining 4,070,208 input points, the vast majority can be evaluated at the default starting precision and do not require precision tuning; we discard these points as well, leaving a total of 567,238 inputs across 469 benchmarks.
@@
All 567,238 are evaluated using both Reval as well as Baseline, a variant of Reval which uses a uniform precision that doubles iteration to iteration. Since Reval and Baseline are the same codebase, just with different tuning strategies, this allows a precision evaluation of Reval’s precision-tuning strategy. To ensure Reval and Baseline are competitive implementations, we also compare both against Sollya, an independently-developed environment for floating-point code development which also uses a uniform precision.202020 Sollya typically doubles the precision in each iteration, but in some cases can detect when it’s “close” to the correct precision and increase the precision by less than doubling. In any case, the precision is uniform. Reval and Baseline consume the benchmark’s standard FPCore formats, but for Sollya we also compile each benchmark into Sollya’s input format; we check that all three tools agree on each input to ensure that the compilation is correct.
@@
For each tool we measure evaluation time and compilation time separately, and focus on evaluation time. Reval and Baseline are called within the same process, so a simple nanosecond timer is used, while for Sollya we use its time command. All three tools are subject to a 20 millisecond timeout per input points. A short timeout is chosen to avoid biasing results due to rare seemingly-infinite loops in Sollya; the timeout is not significant for Reval or Baseline. Evaluation is performed on a machine with an Intel i7-8700 at 3.7 GHz with 32 GB of DDR4 memory running: Ubuntu 22.04; Sollya 7.0 with MPFI 1.5.3, MPFR 4.1.0, and GMP 6.2.1; and Reval 2.0 with Racket 8.11.1. The evaluation code itself is single-threaded, as are Reval and Baseline and, to our knowledge, Sollya, and the evaluation is run without load on the machine.
@@ 7.2. Results
Figure 3 shows evaluations per second for Reval, Baseline, and Sollya, aggregated by Baseline’s chosen precision212121 This is typically also the uniform precision used by Sollya, and typically the largest precision assigned by Reval. and normalized against Baseline’s runtime. Reval is significantly faster on average, by 1.62×1.62$\times$1.62 × compared to Baseline and by 1.47×1.47$\times$1.47 × compared to Sollya. Inputs evaluated at higher-precision are typically harder; Reval is faster than Baseline at any precision and faster than Sollya past $2^{9}$ bits of uniform precision, with Reval’s advantage growing with precision: for the most precision-demanding inputs, Reval is 4.92×4.92$\times$4.92 × faster than Sollya and 2.38×2.38$\times$2.38 × faster than Baseline. In general, Sollya is somewhat faster than Baseline (though note the discussion on timeouts); we suspect this is due to Sollya’s more-careful memory management, orthogonal to Reval’s precision assignment algorithm.
Figure 2. Evaluations per second, with Baseline normalized to 1.0; higher is better. Reval is, on average, 1.62×1.62$\times$1.62 × faster than Baseline and 1.47×1.47$\times$1.47 × faster than Sollya. Inputs are aggregated by Baseline precision, with higher precisions indicating “harder” inputs; Reval has a growing advantage, up to 2.38×2.38$\times$2.38 × over Baseline and 4.92×4.92$\times$4.92 × over Sollya.
Figure 3. A weighted histogram showing the distribution of per-operation execution time by precision, for both Reval and Baseline; lower is better. Baseline wastes a lot of time re-computing the same operations at successively higher precisions, leading to a lot of wasted time at precisions between $2^{6}$ and $2^{11}$. The “tuning” column shows the run time of Reval’s precision tuning algorithm, over all inputs and iterations.
@@
Figure 3 compares per-operation execution time by precision for both Reval and Baseline. Reval spends significantly less time executing operations at precisions between $2^{6}$ and $2^{11}$, because its non-uniform algorithm assigns these high precisions to only those operations that require them. Precision assignment is also fast, as shown by a separate “tuning” column: precision tuning consumes 15.5% of total run time but saves 46.29% of it.
Figure 4. Per-operation precision in final Reval iteration, normalized to maximum precision; left is better. A uniform precision would have all weight at 1.0. Reval instead evaluates most operations at a dramatically lower than maximum precision, with 53.83% of instructions executed at under 20% of the highest precision. This results in its significant performance advantage.
Figure 5. The number of iterations needed for each input in Reval and Baseline; left is better. Reval requires fewer iterations than Baseline because it can derive the correct precision instead of searching for it. This adds an additional minor performance advantage for Reval. The “slack” mechanism is only needed for the second and third iterations, showing it is a minor part of Reval’s algorithm.
Figure 6. The percentage of operations actually evaluated per iteration in Reval; lower is better. Reval’s “repeat optimization” allows it to skip re-evaluating an operation if its precision assignment does not change; Baseline’s uniform precision does not allow this optimization so all its data would be at 100%. 23.77% of operations are skipped on average, adding to Reval’s advantage.
@@
Figure 6 shows the main reason behind Reval’s overall better performance. This figure shows the non-uniform nature of Reval’s precision assignments, with many operations executed at precisions much lower than the maximum precision assigned: 53.83% at less than 20% of the highest precision. By contrast, a uniform precision assignment would have all operations executed at 1.0×1.0$\times$1.0 × the maximum precision. Reval’s ability to assign non-uniform precisions allows it to execute most operations at much lower precisions, which leads to its significant performance advantage due to lower data movement, less computation time, and simpler algorithms.
Figure 6 shows an additional reason behind Reval’s better performance: fewer iterations. A standard uniform precision-doubling algorithm must search for the correct precision by trying higher-and-higher precisions. Reval instead derives the correct precision mathematically, and thus needs fewer search iterations. More precisely, Reval needs only a single tuning iteration for the vast majority (0.73) of inputs, while Baseline needs approximately 3 on average. These extra iterations are wasted computation. Reval’s “slack” mechanism still performs some search, but this still requires far fewer iterations than Baseline because Reval is still deriving the final precision and also because the slack mechanism’s rarer use means it can be tuned more aggressively. For example, after two iterations, 0.24 of Reval’s inputs have converged while only 0.14 of Baseline’s have.
Figure 6 shows a final, minor reason for Reval’s advantage over Baseline: the repeat optimization. Reval’s non-uniform precision assignments often assign the same precisions, iteration to iteration, to an entire subtree, in which case the subtree does not need to be recomputed. Baseline’s uniform precision, meanwhile, requires recomputing every operation at every iteration. On average, Reval skips 23.77% of instructions, with later iterations skipping slightly more.
In summary, non-uniform precisions allow Reval to use much lower precisions than Baseline, evaluates them for fewer iterations, and skips many operation evaluations. None of these advantages are possible with uniform precisions and add up to Reval’s substantial and growing advantage over Baseline. While we weren’t able to gather similarly-granular data about Sollya or another existing real evaluation tool, we expect the same advantages to be present, since all extant tools we know of use the same uniform precision approach.